# Challenges in applying the magnetic field circulation in Ampère-Maxwell’s law

### v1

@@ 6 Conclusion and implication for teaching
In this article, we investigated how students in introductory electromagnetics courses apply magnetic field circulation in the context of the Ampère-Maxwell law. We designed a metacognitive pencil-and-paper question, presented it in writing to 65 students, and then analysed their answers using phenomenography. We complemented our research by conducting interviews with 12 students.
@@
We suggest that in order to promote a deeper understanding of magnetic field circulation in ths context, it is crucial to develop teaching-learning sequences that challenge students to calculate the magnetic field from this law, using unconventional curves and solving inverse problems where the magnetic fields cannot be determined from the line integral [7, 24]. This would allow students to reflect on the relationship between the symmetry of the problem, the fields and the circulation. In addition, it is essential to analyse magnetic field lines in different configurations involving conduction and displacement currents, and to apply the Ampère-Maxwell law to different surfaces bounded by the same curve [25, 26]. It is also worth mentioning the importance of recognising that the Ampère-Maxwell law does not provide information about the sources of the fields, nor does it imply causal relationships, but rather correlations between its different terms [18, 19].
Although the limitations of this study, focusing on a small sample of students, force us to be cautious about its generalisability, we believe that the consistency of the results with other studies on Maxwell’s laws allows us to draw some implications for the teaching of Ampère-Maxwell’s law and magnetic field circulation. In our next studies, we will investigate students’ conceptual difficulties that could arise from cause-and-effect interpretations of Maxwell’s equations, and we will use the results of this work, as well as those of previous research [20, 15], to design a teaching-learning sequence focused on Ampère-Maxwell’s law. This sequence will be implemented in introductory electromagnetism courses and evaluated to determine whether students overcome the conceptual difficulties.

### v2

@@ 6 Conclusion and implication for teaching
In this article we analysed the different qualitative ways in which students in introductory electromagnetics courses apply the magnetic field circulation in the framework of the Ampère-Maxwell law.. We designed a metacognitive pencil-and-paper question, presented it in writing to 65 students, and then analysed their answers using phenomenography. We complemented our research by conducting interviews with 12 students.
@@
Although the small sample of students limits the generalisability of the study, we believe that the consistency of the results with other studies on Maxwell’s laws allows us to draw some implications for the teaching of Ampère-Maxwell’s law and magnetic field circulation. We suggest that in order to promote a deeper understanding of magnetic field circulation in this context and avoid the appearance of reasoning based on functional fixations, it is crucial to develop teaching-learning sequences that challenge students to calculate the magnetic field from this law, using unconventional curves and solving inverse problems where the magnetic fields cannot be determined from the line integral [7, 33]. This would allow students to reflect on the relationship between the symmetry of the problem, the fields and the circulation. In addition, it is essential to analyse magnetic field lines in different configurations involving conduction and displacement currents, and to apply the Ampère-Maxwell law to different surfaces bounded by the same curve [34, 35]. This would help to recognize that the Ampère-Maxwell law does not provide information about the sources of the fields, nor does it imply causal relationships, but rather correlations between its different terms [29, 30]. Finally, non-routinely formulated problems, such as the one presented in this paper, and sense-making activities, such as ‘Ranking Task’ and ‘Conflicting Contentions’, can promote meaningful learning and challenge students’ alternative conceptions. [36, 37], can promote scientifically meaningful learning and challenge students’ alternative conceptions.
In our next studies, we will investigate students’ conceptual difficulties that could arise from cause-and-effect interpretations of Maxwell’s equations, and we will use the results of this work, as well as those of previous research [16, 15], to design a teaching-learning sequence focused on Ampère-Maxwell’s law. This sequence will be implemented in introductory electromagnetism courses and evaluated to determine whether students overcome the conceptual difficulties.
@@ Acknowledgment
Appendix A Protocol of the interview, design and validation
The interview protocol was designed to investigate how students interpret and apply magnetic field circulation in the context of a problem solved by an imaginary student, Agustina. The proposed activity consists of students analysing each step of her solution, in which Agustina chooses a closed semicircular curve to calculate the magnetic field associated with a varying electric field. Working in groups of two or three, students have to verbalise their reasoning aloud during semi-structured interviews and reflect on each step of the resolution. This approach facilitates an exploration of students‘ conceptual difficulties, in line with the phenomenographic approach to capture variations in students’ understanding.
The protocol is organised in three phases to allow for a continuous and detailed exploration of the students’ ideas and reasoning.
1. Introduction and initial understanding phase. The aim of this phase is to assess initial understanding of the problem, to identify concepts that students consider relevant and to observe how they perceive the phenomenon collaboratively. Students are presented with the problem situation: a variable electric field confined to a cylindrical region of radius R and Agustina’s intention to determine the magnetic field at a point located at a distance a from the centre of the region. Opening question: ‘Please have one of you read out the description of the question. Explain in your own words what you think is being asked in the activity’.
2. Clarification and reflection phase on the proposed resolution. The aim of this phase is to capture how students experience and understand the problem, identifying conceptual difficulties and ways of reasoning in each step of the resolution. Students analyse as a group each step followed by Agustina in her resolution. Reflection is encouraged through non-directive questions and follow-up questions, such as: ‘What do you mean by that?’, ‘Do you agree with this step? Why?’, ‘What would you change in this solution?’, ‘What do you think about the choice of curve C to calculate the magnetic field?’, ‘What do you think about the way the magnetic field circulation is calculated?’. Students discuss their agreements and disagreements and suggest changes if they think any adjustments to the proposed resolution are necessary.
3. Justification and conclusions phase. The aim of this phase is to stimulate a comprehensive and shared understanding of the problem, allowing students to express any additional reasoning and reflect collectively on the solution. Students are invited to formulate general conclusions about the steps followed by Agustina, justifying their opinions and consolidating their ideas. Final question: ‘After analysing the steps followed by Agustina, what do you think as a group about the procedure she used to solve the problem, would you change any aspect or do you think the process correctly reflects what should be done to calculate the magnetic field in this case?
The design and validation of the protocol was conducted iteratively, while integrating elements of initial validation into design decisions. The initial draft was reviewed by two PER experts to assess the relevance, clarity and ability of the protocol to capture variations in student experiences. This feedback helped to ensure that the questions promoted collaborative analysis without influencing students’ reasoning by adjusting the language and clarifying the questions. Final validation was carried out through a pilot test with two students. The results allowed refining the structure of the protocol, eliminating possible ambiguities and ensuring that both the non-directive and follow-up questions were clear, open-ended and did not interfere with students’ collaborative reasoning. This process ensured that the protocol allowed for a wide range of interpretations and reasoning around the application of magnetic field circulation to be observed, aligning with the phenomenographic approach of the study.